OpenStudy (anonymous):

18 x 3 \[\div\] 3^3 ?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Well 18*3 is just 54. 3^3=(3*3*3)=(9*3)=27. So 54/27=2

OpenStudy (anonymous):

Quicker than that: (18*3)/(3^3)=18/3^2=18/9=2

OpenStudy (anonymous):

18 x 3 = 2 X 3^3 SO DIVISION BY 3^3 ... ...is 2